# SoundMorpher: Perceptually-Uniform Sound Morphing with Diffusion Model

### v1

@@ Abstract
We present SoundMorpher, a sound morphing method that generates perceptually uniform morphing trajectories using a diffusion model. Traditional sound morphing methods models the intractable relationship between morph factor and perception of the stimuli for resulting sounds under a linear assumption, which oversimplifies the complex nature of sound perception and limits their morph quality. In contrast, SoundMorpher explores an explicit proportional mapping between the morph factor and the perceptual stimuli of morphed sounds based on Mel-spectrogram. This approach enables smoother transitions between intermediate sounds and ensures perceptually consistent transformations, which can be easily extended to diverse sound morphing tasks. Furthermore, we present a set of quantitative metrics to comprehensively assess sound morphing systems based on three objective criteria, namely, correspondence, perceptual intermediateness, and smoothness. We provide extensive experiments to demonstrate the effectiveness and versatility of SoundMorpher in real-world scenarios, highlighting its potential impact on various applications such as creative music composition, film post-production and interactive audio technologies. See SoundMorpher-demo for the listening demonstration.
1 Introduction
Sound morphing is a technique to create a seamless transformation between multiple sound recordings. The goal is to produce perceptual intermediate sounds that gradually change from one sound to another. Sound morphing has a wide range of applications, including music compositions, synthesizers, psychoacoustic experiments to study timbre spaces (Caetano & Rodet, 2011; Hyrkas, 2021), and practical applications such as film post-production, AR or VR interactive games, and adaptive audio content in video games (Qamar et al., 2020; Siddiq, 2015).
@@ 1 Introduction
Traditional sound morphing methods used the interpolation principle of sound synthesis technique, which relies on interpolating the parameters of a sinusoidal model (Tellman et al., 1995; Osaka, 1995; Williams et al., 2014). Others make use of digital signal processing techniques to explore high-level audio features in the time-frequency domain to achieve more effective and continuous morphing (Williams et al., 2014; Brookes & Williams, 2010; Caetano & Rodet, 2010; 2011; Roma et al., 2020; Caetano, 2019). However, these methods are limited to applications such as producing inharmonic and noisy environmental sounds (Gupta et al., 2023; Kamath et al., 2024). Despite the increasing interest in applying machine learning to sound generation, there has only been limited exploration in sound morphing. Recent approaches (Zou et al., 2021; Gupta et al., 2023; Kim et al., 2019b; Kamath et al., 2024) have shown their superior effectiveness compared to traditional methods in various scenarios. However, we observed several critical limitations of those existing methods. Firstly, they are primarily designed for static or cyclostationary morphing (see Sec. 3.1), limiting their applicability to dynamic sound transformations. Secondly, these approaches often lack sufficient quantitative evaluation, limiting further analysis of their effectiveness. Thirdly, they require training on task-specific datasets, which limits their application in different scenarios. Most importantly, they typically assume a linear relationship between morphing factors and sound perception, and achieve smooth morphing by gradually changing the morph factor. This assumption oversimplifies the complex nature of sound perception, as gradually changing morph factors does not inherently result in smooth perceptual transitions. To this end, our goal is to develop a method that achieves perceptually coherent morphing, ensuring seamless and natural sound transition.
In this paper, we introduce SoundMorpher, a sound morphing method that produces perceptually smooth and intermediate morphing, comprising the following key contributions.
@@
  • SoundMopher is the first open-world sound morphing method based on a pre-trained diffusion model, which integrates typical morph tasks such as static, dynamic and cyclostationary morphing. Unlike prior works (Kim et al., 2019b; Gupta et al., 2023), SoundMorpher can be broadly applied to various real-world tasks without requiring extensive retraining.
@@
  • We propose the sound perceptual distance proportion (SPDP), which explicitly connects morph factors and perceptual stimuli of morphed results. This allows SoundMorpher to produce morphing paths with a uniform change in perceptual stimuli, achieving more seamless perceptual transitions compared to existing methods (Kamath et al., 2024).
@@
  • We adapt a set of comprehensive quantitative metrics according to criteria proposed by Caetano & Osaka (2012) for evaluation, addressing the lack of quantitative assessment for sound morphing systems (Caetano, 2019; Zou et al., 2021; Caetano & Rodet, 2013) and may offer insights for analyzing and comparing future sound morphing methods.
@@
  • We provide extensive experiments to demonstrate that SoundMorpher can be effectively applied to several potential applications in broader real-world scenarios, including musical instrument timbre morphing, music morphing and environmental sound morphing.
2 Related Work
@@
In this section, we first present a detailed review of related works on sound morphing task. Then, we also briefly introduce tasks that are similar to sound morphing and clarify the differences.
@@
Sound morphing. Traditional sound morphing methods rely on interpolating parameters of a sinusoidal sound synthesis model (Tellman et al., 1995; Osaka, 1995; Williams et al., 2014; Primavera et al., 2012). To achieve more effective and continuous morphing, Williams et al. (2014); Brookes & Williams (2010); Caetano & Rodet (2010; 2011); Roma et al. (2020); Caetano (2011) target on exploring perceptual spectral domain audio features by digital signal processing techniques, such as MFCCs, spectral envelope, etc.. Others such as Kazazis et al. (2016) involve a hybrid approach that extracts audio descriptors to morph accordingly and interpolate between the spectrotemporal fine structures of two endpoints according to morph factors. Machine learning sound morphing methods offer advantages such as high morphing quality by leveraging semantic representation interpolation within a model instead of traditional audio feature interpolation. Zou et al. (2021) proposes a non-parallel many-to-one static timbre morphing framework that integrates and fine-tunes the machine learning technique (i.e., DDSP-autoencoder (Engel et al., 2020)) with spectral feature interpolation (Caetano & Rodet, 2013). Kim et al. (2019b) targets synthesizing music corresponding to a note sequence and timbre, which uses non-linear instrument embedding as timbre control parameters under a pretrained WaveNet (Engel et al., 2017) to achieve timbre morphing between instruments. Luo et al. (2019) learns latent distributions of VAEs to disentangle representations for pitch and timbre of musical instrument sounds. Tan et al. (2020) uses a GM-VAE to achieve style morphing to generate realistic piano performances in the audio domain following temporal style conditions for piano performances, which morphs the conditions such as onset roll and MIDI note into input audio. MorphGAN (Gupta et al., 2023) targets on audio texture morphing by interpolating within conditional parameters, and trained the model on a water-wind texture dataset. A recent concurrent work by Kamath et al. (2024) uses a pre-trained AudioLDM (Liu et al., 2023) to morph sound between two text prompts. In contrast, we focus on classical sound morphing, where the morphing process is performed directly between two given audios rather than between text prompts. A key advantage of our method is its ability to provide precise guidance during the morphing process, as the target audio delivers exact information on how the source sound should evolve—something that text prompts cannot always achieve, for example, morphing between two music compositions.
@@
Synthesizer preset interpolation. Synthesizer preset interpolation achieves sound morphing by developing models that compute interpolations within the domain of synthesis parameters for a black-box synthesizer (Le Vaillant & Dutoit, 2023; Dutoit et al., 2023; Le Vaillant & Dutoit, 2024). Unlike classical sound morphing, which perceptually blends two audio files into an intermediate sound, synthesizer preset interpolation treats the synthesizer as a non-differentiable black box, with presets composed of both numerical and categorical parameters. By smoothly interpolating between these presets, the task aims to achieve seamless morphing of synthesized sounds.
@@
Text-to-audio editing. Text-to-audio editing is the process of using text queries to edit audio. With the success of diffusion models in image editing tasks, recent works target zero-shot audio editing with text instructions (Manor & Michaeli, 2024; Zhang et al., 2024; Lan et al., 2024) involving tasks such as inpainting, outpainting, timbre transfer, music genre transfer, or vocals removal.
@@
Timbre transfer. Timbre transfer is a specific task that aims at converting the sound of a musical piece by one instrument (source) into the same piece played by another instrument (target). This concerns the task of converting a musical piece from one timbre to another while preserving the other music-related characteristics (Comanducci et al., 2024; Jain et al., 2020; Li et al., 2024).
@@
Voice conversion and morphing. Voice conversion (VC) involves modifying vocal characteristics of a source speech to match a target speaker, either by using target speeches or text (Li et al., 2023; Yao et al., 2024; Niu et al., 2024; Sheng et al., 2024). The primary objective of VC is to alter the vocal identity to closely resemble the target voice style, while preserving the linguistic content of the source speech. Voice morphing is a broader scope, focusing on blending or transforming one voice into another. This often involves creating an intermediate voice that incorporates characteristics of both source and target voices, allowing for gradual transitions between them (Sheng et al., 2024).
3 Preliminaries
3.1 Sound Morphing
@@
Sound morphing aims to produce intermediate sounds as different combinations of model source sound $\hat{S}_{1}$ and target sound $\hat{S}_{2}$ (Caetano & Rodet, 2010; 2011), which can be formulated as
@@
Each step is characterized by one value of a single parameter $\alpha$, the so-called morph factor, which ranges between 00 and $1$, where $\alpha=0$ and $\alpha=1$ produce resynthesized source and target sounds, respectively. Due to the intrinsic temporal nature of sounds, sound morphing usually involve three main types: dynamic morphing, where $\alpha$ gradually transfers from 00 to $1$ over time dimension (Kazazis et al., 2016), static morphing, where a single morph factor $\alpha$ leads to an intermediate sound between source and target (Sethares & Bucklew, 2015), and cyclostationary morphing where several hybrid sounds are produced in different intermediate points (Slaney et al., 1996).
@@
To solve the limitation on previous works that target on expensive perceptual evaluation only, Caetano & Osaka (2012) proposes three objective criteria for sound morphing techniques: (1) Correspondence. The morph is achieved by a description whose elements are intermediate between source and target sounds, highlighting semantic level transition; (2) Intermediateness. The morphed objects should be perceived as intermediate between source and target sounds, evaluating perceptual level correlation; (3) Smoothness. The morphed sounds should change gradually (i.e., ‘smoothly’) from source to target sounds, by the same amount of perception increment. Under the assumption of linear perceptual stimuli, adding the same factor should increase the same amount of perception. In this study, we evaluate SoundMorpher according to the three criteria by a series of comprehensive objective quantitative metrics in addition to perceptual evaluation.
3.2 Latent Diffusion Model on Audio Generation
@@
SoundMorpher utilizes a pretrained text-to-audio (TTA) latent diffusion model (LDM) (Rombach et al., 2022) to achieve sound morphing. This approach offers the advantage of performing various types of sound morphing without the need to train the entire model or use additional datasets. Specifically, we use AudioLDM2 (Liu et al., 2024), a multi-modality conditions to audio model. It employs a pre-trained variational autoencoder (VAE) (Kingma & Welling, 2013) to compress audio $x$ into a low-dimension latent space as VAE representations $z$. AudioLDM2 generates latent variables $z_{0}$ from a Gaussian noise $z_{T}$ given the condition $C$ and further reconstruct audio $\hat{x}$ from $z_{0}$ by VAE decoder and a vocoder (Kong et al., 2020). AudioLDM2 uses an intermediate feature $Y$ as an abstraction of audio data $x$ to bridge the gap between conditions $C$ and audio $x$, named language of audio (LOA). The LOA feature is obtained by a AudioMAE (Huang et al., 2022; Tan et al., 2024) and a series of post-processing formulated as $Y=\mathcal{A}(x)$. The generation function $\mathcal{G}(\cdot)$ is achieved by a LDM. In the inference phase, AudioLDM2 approximates LOA feature by the given condition as $\hat{Y}=\mathcal{M}(C)$ using a fine-tuned GPT-2 model (Radford et al., 2019). Then generates audios conditioned on the estimated LOA feature $\hat{Y}$ and an extra text embedding $E_{T5}$ from a FLAN-T5 (Chung et al., 2024) with a LDM as $\hat{x}=\mathcal{G}(\hat{Y},E_{T5})$. We denote the conditional embeddings in AudioLDM2 as $E=\{\hat{Y},E_{T5}\}$, therefore, the generative process becomes $\hat{x}=\mathcal{G}(E)$.
@@
Diffusion Models. The LDM performs a forward diffusion process during training, which is defined as a Markov chain that gradually adds noise to the VAE representation $z_{0}$ over $T$ steps as $z_{t}=\sqrt{1-\beta_{t}}z_{t-1}+\sqrt{\beta_{t}}\epsilon_{t}$. where $\epsilon_{t}\sim N(0,I)$ and noise schedule hyperparameter $\beta_{t}\in[0,1]$. Therefore, we can derive the distribution of $z_{t}$ given $z_{0}$ as $q(z_{t}|z_{0})=\sqrt{\gamma_{t}}z_{0}+\sqrt{1-\gamma_{t}}\epsilon_{t}$, where $\gamma_{t}=\prod_{t=1}^{t}1-\beta_{t}$. The LDM learns a backward transition $\epsilon_{\theta}(z_{t},t)$ from the prior distribution $N(0,I)$ to the data distribution $z$, that predicts the added noise $\epsilon_{t}$ (Ho et al., 2020). Following the objective function of denoising diffusion probabilistic models (DDPM) (Ho et al., 2020), the objective function for training AudioLDM2 is
$\text{min}_{\theta}\mathcal{L}_{DPM}=\text{argmin}_{\theta}[\mathbb{E}_{z_{0},E,t\sim\{1,...,T\}}||\epsilon_{\theta}(z_{t},E,t)-\epsilon_{t}||_{2}^{2}]$ (2)
@@
To reduce computational demands on inference, AudioLDM2 uses denoising diffusion explicit models (DDIM) (Song et al., 2020), which provides an alternative solution and enables significantly reduced sampling steps with high generation quality. The DDIM reverse diffusion process is
$z_{t-1}=\sqrt{\gamma_{t-1}}(\frac{z_{t}-\sqrt{1-\gamma_{t}}\epsilon_{\theta}(z_{t},E,t))}{\sqrt{\gamma_{t}}})+\sqrt{1-\gamma_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}(z_{t},E,t)+\sigma_{t}\epsilon_{t}$ (3)
We can revise a deterministic mapping between $z_{0}$ and its latent state $z_{T}$ once the model is trained (Dhariwal & Nichol, 2021; Yang et al., 2023) by the following equation
@@
4 Method

### v2

@@ Abstract
We present SoundMorpher, an open-world sound morphing method designed to generate perceptually uniform morphing trajectories. Traditional sound morphing techniques typically assume a linear relationship between the morphing factor and sound perception, achieving smooth transitions by linearly interpolating the semantic features of source and target sounds while gradually adjusting the morphing factor. However, these methods oversimplify the complexities of sound perception, resulting in limitations in morphing quality. In contrast, SoundMorpher explores an explicit relationship between the morphing factor and the perception of morphed sounds, leveraging log Mel-spectrogram features. This approach further refines the morphing sequence by ensuring a constant target perceptual difference for each transition and determining the corresponding morphing factors using binary search. To address the lack of a formal quantitative evaluation framework for sound morphing, we propose a set of metrics based on three established objective criteria. These metrics enable comprehensive assessment of morphed results and facilitate direct comparisons between methods, fostering advancements in sound morphing research. Extensive experiments demonstrate the effectiveness and versatility of SoundMorpher in real-world scenarios, showcasing its potential in applications such as creative music composition, film post-production, and interactive audio technologies.
Index Terms:
Sound morphing, open-world, diffusion models.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
Sound morphing is a technique to create a seamless transformation between multiple sound recordings. The goal is to produce intermediate sounds that might exist between two sounds and contrasts with simpler mixing or cross-fading between sounds. Sound morphing has a wide range of applications, including music compositions, synthesizers, psychoacoustic experiments to study timbre spaces [1, 2], film post-production, AR/VR, and adaptive audio content in video games [3, 4]. Traditionally, sound morphing relied on interpolating the parameters of a sinusoidal synthesis model [5, 6, 7]. High-level audio features in the time-frequency domain can also be controlled to achieve more effective and continuous morphing [7, 8, 9, 1, 10, 11]. These techniques have quality limitations that rule out applications in inharmonic and noisy sounds such as environmental sounds [12, 13]. Increasing interest in sound generation using machine learning (ML) has delivered techniques for ML-based sound morphing [14, 12, 15, 13] that out-perform traditional methods in scenarios such as timbre morphing, audio texture morphing and environmental sound morphing; however, we observed critical limitations of these existing sound morphing methods. Firstly, they are primarily designed for specific morphing methods or tailored to particular application scenarios on task-specific datasets, which limits their ability to broader applications and different scenarios. Secondly, a lack of sufficient and formal quantitative evaluation limits further analysis of their effectiveness [12] and impedes fair comparisons with other sound morphing techniques. Most importantly, these methods typically assume a linear relationship between morphing factors and sound perception, as proposed by [16]. They achieve smooth morphing by interpolating semantic features linearly between the source and target sounds and gradually varying the morphing factor. This assumption oversimplifies the complex nature of sound perception, as gradually changing morph factors does not inherently result in smooth perceptual transitions. Our goal is to address these limitations by introducing SoundMorpher, a flexible sound morphing method, which can be broadly applied to different sound morphing tasks to achieve perceptually coherent morph, ensuring seamless and natural sound transition. In this work we:
@@
  • Introduce the first open-world sound morphing method based on a pre-trained diffusion model, which integrates typical morph tasks such as static, dynamic and cyclostationary morphing. Unlike [15, 12], SoundMorpher can be broadly applied to various real-world sound morphing tasks without requiring extensive retraining on an additional dataset, paving the way for future sound morphing studies to focus on leveraging pre-trained audio generation models for greater efficiency and versatility.
@@
  • Propose sound perceptual distance proportion (SPDP), which explicitly connects morph factors and perception of morphed results. This allows SoundMorpher to produce morphing paths with a uniform change in perceptual stimuli, achieving more seamless perceptual transitions compared to existing methods [13].
@@
  • Adapt established criteria [16] for quantitative evaluation, addressing the lack of comprehensive objective quantitative assessment for sound morphing systems [11, 14, 17].
@@
  • Provide extensive experiments to demonstrate that SoundMorpher can be effectively applied to various applications in real-world scenarios including timbre morphing, music morphing and environmental sound morphing.
II Related Work
@@
In this section, we present a detailed review of related works on sound morphing task and briefly introduce contrasts with other similar tasks.
@@
Sound morphing. Traditional sound morphing methods are based on interpolating parameters of a sinusoidal sound synthesis model [5, 6, 7, 18]. To achieve more effective and continuous morphing, [7, 8, 9, 1, 10, 19] exploring perceptual spectral domain audio features by digital signal processing techniques, such as MFCCs, spectral envelope, etc. A hybrid approach has been explored that extracts audio descriptors to morph accordingly and interpolates between the spectrotemporal fine structures of two endpoints according to morph factors [20]. Machine learning sound morphing methods offer advantages such as high morphing quality by interpolation of semantic representations within a model instead of traditional audio features. [14] proposes a non-parallel many-to-one static timbre morphing framework that integrates and fine-tunes the machine learning technique (i.e., DDSP-autoencoder [21]) with spectral feature interpolation [17]. [15] synthesizes music corresponding to a note sequence and timbre, using nonlinear instrument embeddings as timbre control parameters under a pretrained WaveNet [22] to achieve timbre morphing between instruments. [23] learns latent distributions of VAEs to disentangle representations of the pitch and timbre of musical instrument sounds. [24] uses a GM-VAE to achieve style morphing to generate realistic piano performances in the audio domain following temporal style conditions for piano performances, which morphs the conditions such as onset roll and MIDI note into input audio. MorphGAN [12] targets audio texture morphing by interpolating within conditional parameters training the model on a water-wind texture dataset. A recent concurrent work [13] uses a pre-trained AudioLDM [25] to morph sound between two text prompts. In contrast, we focus on classical sound morphing, where the morphing process is performed directly between two given sounds rather than between text prompts. A key advantage of our method is its ability to provide precise guidance during the morphing process, since the target audio delivers exact information on how the source sound should evolve—something that text prompts cannot always achieve, for example, morphing between two music compositions.
@@
Synthesizer preset interpolation. A form of sound morphing can be achieved by developing models that compute interpolations in the parameters for a black-box synthesizer [26, 27, 28]. Unlike classical sound morphing, which perceptually blends two audio files into an intermediate sound, synthesizer preset interpolation treats the synthesizer as a non-differentiable black box, with presets composed of both numerical and categorical parameters. By smoothly interpolating between these presets, the task aims to achieve seamless morphing of synthesized sounds.
@@
Text-to-audio editing. Text-to-audio editing is the process of using text queries to edit audio. With the success of diffusion models in image editing tasks, recent works target zero-shot audio editing with text instructions [29, 30, 31] involving tasks such as inpainting, outpainting, timbre transfer, music genre transfer, or vocals removal.
@@
Timbre transfer. Timbre transfer is a specific task that aims at converting the sound of a musical piece by one instrument (source) into the same piece played by another instrument (target). This concerns the task of converting a musical piece from one timbre to another while preserving the other music-related characteristics [32, 33, 34].
@@
Voice conversion and morphing. Voice conversion (VC) involves modifying vocal characteristics of a source speech to match a target speaker, either by using target speeches or text [35, 36, 37, 38]. The primary objective of VC is to alter the vocal identity to closely resemble the target voice style while preserving the linguistic content of the source speech. Voice morphing is a broader scope, focusing on blending or transforming one voice into another. This often involves creating an intermediate voice that incorporates characteristics of both source and target voices, allowing for gradual transitions between them [38].
III Preliminaries
III-A Sound Morphing
@@
Sound morphing aims to produce intermediate sounds as different combinations of model source sound $\hat{S}_{1}$ and target sound $\hat{S}_{2}$ [9, 1], which can be formulated as
@@
Each step is characterized by one value of a single parameter $\alpha$, the so-called morph factor, which ranges between 00 and $1$. $\alpha=0$ and $\alpha=1$ produce resynthesized source and target sounds, respectively. Due to the intrinsic temporal nature of sounds, sound morphing usually involves three main types: dynamic morphing, where $\alpha$ gradually transfers from 00 to $1$ over time [20], static morphing, where a single morph factor $\alpha$ leads to an intermediate sound between source and target [39], and cyclostationary morphing where several hybrid sounds are produced in different intermediate points [40].
@@
To solve the limitation on previous works that target on expensive perceptual evaluation only, [16] provided three objective criterions to formalize the evaluation for sound morphing methods: (1) Correspondence. The morph is achieved by a description whose elements are intermediate between source and target sounds, highlighting semantic level transition; (2) Intermediateness. The morphed objects should be perceived as intermediate between source and target sounds, evaluating perceptual level correlation; (3) Smoothness. The morphed sounds should change gradually from source to target sounds, by the same amount of perception increment. Specifically, the linear assumption proposed by [16] is that adding the same amount of morph factor should increase the same amount of perception difference.
In this study, we evaluate SoundMorpher according to the three criterions proposed by [16] with a series of comprehensive objective quantitative metrics.
III-B Latent Diffusion Model on Audio Generation
@@
SoundMorpher uses a pretrained text-to-audio (TTA) latent diffusion model (LDM) [41] to achieve sound morphing. This approach offers the advantage of performing various types of sound morphing without the need to train the entire model or use additional datasets. Specifically, we use AudioLDM2 [42], a multi-modality conditions to audio model. It employs a pre-trained variational autoencoder (VAE) [43] to compress audio $x$ into a low-dimension latent space as VAE representations $z$. AudioLDM2 generates latent variables $z_{0}$ from a Gaussian noise $z_{T}$ given the condition $C$ and further reconstructs audio $\hat{x}$ from $z_{0}$ by VAE decoder and a vocoder [44]. AudioLDM2 uses an intermediate feature $Y$ as an abstraction of audio data $x$ to bridge the gap between conditions $C$ and audio $x$, named language of audio (LOA). The LOA feature is obtained by a AudioMAE [45, 46] and a series of post-processing formulated as $Y=\mathcal{A}(x)$. The generation function $\mathcal{G}(\cdot)$ is achieved by a LDM. In the inference phase, AudioLDM2 approximates LOA feature by the given condition as $\hat{Y}=\mathcal{M}(C)$ using a fine-tuned GPT-2 model [47]. Then generates audios conditioned on the estimated LOA feature $\hat{Y}$ and an extra text embedding $E_{T5}$ from a FLAN-T5 [48] with a LDM as $\hat{x}=\mathcal{G}(\hat{Y},E_{T5})$. We denote the conditional embeddings in AudioLDM2 as $E=\{\hat{Y},E_{T5}\}$, therefore, the generative process becomes $\hat{x}=\mathcal{G}(E)$.
@@
Diffusion Models. The LDM performs a forward diffusion process during training, which is defined as a Markov chain that gradually adds noise to the VAE representation $z_{0}$ over $T$ steps as $z_{t}=\sqrt{1-\beta_{t}}z_{t-1}+\sqrt{\beta_{t}}\epsilon_{t}$. where $\epsilon_{t}\sim N(0,I)$ and noise schedule hyperparameter $\beta_{t}\in[0,1]$. Therefore, we can derive the distribution of $z_{t}$ given $z_{0}$ as $q(z_{t}|z_{0})=\sqrt{\gamma_{t}}z_{0}+\sqrt{1-\gamma_{t}}\epsilon_{t}$, where $\gamma_{t}=\prod_{t=1}^{t}1-\beta_{t}$. The LDM learns a backward transition $\epsilon_{\theta}(z_{t},t)$ from the prior distribution $N(0,I)$ to the data distribution $z$, that predicts the added noise $\epsilon_{t}$ [49]. Following the objective function of denoising diffusion probabilistic models (DDPM) [49], the objective function for training AudioLDM2 is
$\min\limits_{\theta}\mathcal{L}_{\text{DPM}}=\min_{\theta}[\mathbb{E}_{z_{0},E,t\sim\{1,..,T\}}||\epsilon_{\theta}(z_{t},E,t)-\epsilon_{t}||_{2}^{2}]$ (2)
@@
To reduce computational demands on inference, AudioLDM2 uses denoising diffusion explicit models (DDIM) [50], which provides an alternative solution and enables significantly reduced sampling steps with high generation quality. We can revise a deterministic mapping between $z_{0}$ and its latent STATE $z_{T}$ once the model is trained [51, 52] by the equation
@@
Figure 1: Overview of SoundMorpher pipeline, where the snowflake represents the model parameters are frozen.
IV Method